# PlanScope: Learning to Plan Within Decision Scope Does Matter

### v1

@@ Abstract
In the context of autonomous driving, learning-based methods have been promising for the development of planning modules. During the training process of planning modules, directly minimizing the discrepancy between expert-driving logs and planning output is widely deployed. In general, driving logs consist of suddenly appearing obstacles or swiftly changing traffic signals, which typically necessitate swift and nuanced adjustments in driving maneuvers. Concurrently, future trajectories of the vehicles exhibit their long-term decisions, such as adhering to a reference lane or circumventing stationary obstacles. Due to the unpredictable influence of future events in driving logs, reasoning bias could be naturally introduced to learning based planning modules, which leads to a possible degradation of driving performance. To address this issue, we identify the decisions and their corresponding time horizons, and characterize a so-called decision scope by retaining decisions within derivable horizons only, to mitigate the effect of irrational behaviors caused by unpredictable events. This framework employs wavelet transformation based log preprocessing with an effective loss computation approach, rendering the planning model only sensitive to valuable decisions at the current state. Since frequency domain characteristics are extracted in conjunction with time domain features by wavelets, decision information across various frequency bands within the corresponding time horizon can be suitably captured. Furthermore, to achieve valuable decision learning, this framework leverages a transformer based decoder that incrementally generates the detailed profiles of future decisions over multiple steps. Our experiments demonstrate that our proposed method outperforms baselines in terms of driving scores with closed-loop evaluations on the nuPlan dataset. The source code and associated videos are available at https://github.com/Rex-sys-hk/PlanScope.
@@ I Introduction
Learning-based autonomous driving systems have been a promising approach for the popularity of unmanned vehicles. In particular, learning from driving logs of experts is broadly adopted to supervise neural networks for the planning task. During the training process, distance error is commonly employed, which is to measure the discrepancy between the planned state sequence and the expert demonstrations, thereby optimizing parameters of the neural network [1]. Yet, when training end-to-end autonomous driving systems, it is not entirely reasonable to set the logged expert states as target sequence, because future events are inherently uncertain, and potential disruptions such as unexpected obstacles or fluctuating traffic signals can significantly alter the trajectory planning. These unpredictable elements necessitate agile and nuanced adjustments to driving behavior. Some seemingly predictable decisions, such as the interaction behavior and future trajectory of other traffic participants, are difficult to accurately predict. For example, when modeling the interactions by non-cooperative game theory, the decision planner converges only when observing other traffic participants respond. Otherwise, one can only avoid interactions with others to evade falling into disadvantageous states. Meanwhile, a short-sighted planner is also not desirable, as the planner also needs to keep its long-term decision in consideration when choosing sudden maneuvers. Essentially, the state sequence of experts recorded in the log is a combination of short-term and long-term decisions under different scopes. A diagram of this concept is shown in Fig. 1.
@@
Obviously, an agent needs a broad enough view of the overall situation, handles the current trivialities, and does not overly concern about the trivialities of the future. As a reference, in traditional planning methods, the process is artificially segmented into distinct tiers according to the problem scope: path planning, behavior planning, and motion planning, with each addressed by specialized modules [2]. Conversely, learning-from-demonstration approaches typically involve directly learning future trajectories of a predetermined length, because the nuanced adjustment components and long-term decision components cannot be directly distinguished from log replay. Driving-behavior-reasoning by manual annotation could be an approach to analyze the long-term and short-term decisions [3], by which the behavior duration and their starting time can be obtained. However, this paradigm requires a heavy workload and is typically challenging to scale up. To address the aforementioned challenges, we employ wavelet transformation to extract information at various frequency bands, enabling our model to discern the appropriate actions from log replays in a self-supervised manner. Unlike the Fourier transform, wavelet transformation utilizes basis functions of finite length, which facilitates the capture of choppy features in signals [4]. This technique employs a convolution-like process to extract signal components. Based on wavelet transformation, we propose a mechanism to learn within the decision scope, such that the time horizon of distinct frequency bands can be modulated. In this sense, we can compare the decision details generated from the neural network with decomposed details at different horizons, and this prevents the model from overextending its planning into the future. In order to render the details obtained by decomposition efficiently in model training, we introduce a decoding module that generates detailed short-term decisions at multiple detail levels recursively. Our contributions are threefold:
  • We introduce a novel self-supervised mechanism that learns to plan within decision scope. This approach leverages wavelet decomposition, which extracts frequency domain characteristics in conjunction with time domain features, to capture motion information across various scales within the corresponding time horizon.
  • We propose a new transformer based Iterative Detail Decoder (IDD) to iteratively generate the trajectory decisions and its finer details, thereby enabling models to selectively learn the trajectory intricacies.
@@
  • Our method has been rigorously evaluated on the nuPlan [5] dataset, and the results demonstrate that it achieves superior driving scores compared to baseline methods in closed-loop evaluations, especially under critical safety related metrics.
@@ II-A Horizon Related Planner Design
Sequential decision-making can be encapsulated with the framework of Partially Observable Markov Decision Processes (POMDPs) [6]. Within this paradigm, the ego agent must make decisions based on environmental observations, anticipate subsequent states following its actions, and maximize the cumulative reward. Given the world model (also known as the transition function) and the constraint of limited observations, the optimal policy can be derived by exploring the state space and resolving the Bellman equation. However, in autonomous driving scenarios, the world model is generally unknown, making it challenging to determine the optimal policy sequence incrementally in real-time [7]. Hence, under traditional planning frameworks, hierarchical structure is typically employed to streamline computations [8, 9]. In the hierarchical structure, the planning problem is decomposed across various frequency bands, each with distinct hierarchical tendencies. For large-scale exploration, one might opt for a lower resolution and focus on non-real-time targets. Conversely, within the smaller scale areas identified by the large-scale search, our planner can leverage higher resolution and account for dynamic objects that are impractical to predict accurately. In [10], it adapts this concept to the prediction phase, and segments the forecast for the subsequent four seconds into two stages. Initially, it meticulously predicts the agent’s reachable area at 0.1-second intervals for the first 0.5 seconds. Subsequently, for the remaining 3.5 seconds, it utilizes an adaptive filter to depict potential vehicle positions. This approach allows the agent to maintain safety at a fine scale while preserving adaptability at a broader scale. In terms of learning-based planning, analogous strategies can be harnessed to simplify calculations and enhance performance. For instance, the Negative Log Likelihood (NLL) Loss, which is frequently utilized in predictive tasks [11], assigns greater uncertainty to long-term trajectory points, such that the scarcity of future information can be reflected. This phenomenon of probabilistic regression result helps to posit that historical trajectories provide limited insight, sufficient only for fine-tuned operations in the short term and for general directional guidance in the long term. However, NLL loss or probabilistic predictions are incompatible with the planning module, which necessitates deterministic outputs for trajectory followers. It leaves an open and interesting problem to develop an approach to address this trajectory planning challenge.
@@ III Methodology
Figure 2: The overall model framework initiates with the individual encoding of driving context elements, including the Ego Vehicle, Dynamic Agents, Static Objects, High-Definition Map, Traffic Signs, and Signals, followed by their concatenation with positional encoding (PE) [17]. This concatenated context is subsequently processed through a Multi-layer Transformer encoder, iterated $L_{enc}$ times. The scenario encoding is then decoded to formulate agent predictions, which serve to comprehensively supervise the model’s understanding of the driving context. The reference lane encoding and anchor-free variables are combined to form a query for the scenario encoding. The decoder’s output generates residual latent variables, with each layer’s latent variable being logged as that layer’s detail and aggregated with the reference query to form the new query for the subsequent layer.
@@
where $f$ denotes the neural network model, $\theta$ is the model parameter, $\mathcal{L}$ is the loss function, and $\hat{G}_{t}$ is the ground truth of $G_{t}$ from logged context.
@@
where $\mathcal{L}_{\text{pre}}$ calculates the future trajectory discrepancy between generated and that from dataset, $\mathcal{L}_{\text{col}}$ computes the invasion distance using $N_{c}$ circle represented vehicle outline in future $T_{f}$ steps as described in [21], $R_{c}$ is the summed radius of ego and agents outline circle, $\epsilon$ is the tolerance, and $d_{i}^{t}$ represents the center distance between outline circles at time step $t$. Drawing inspiration from [22, 23, 24], the decoder’s query is formulated as a concatenation of the reference lane encoding and anchor-free variables. The reference lane encoded features are concatenated with anchor-free learnable queries and serve as the initial query $Q_{0}$ to the scenario encoding within multi-layer self-attention operations, facilitating the generation of residual quantities of $Q$ for subsequent layers. More details on the acquisition of supervision signals and the decoding of detailed information are elaborated in Section III-A and Section III-B, respectively.
III-A Expert Log Decomposition
Figure 3: Velocity along the trajectory is decomposed into different levels of detail. Only the red cross-marked data points in each level are in scope.
@@ III-A Expert Log Decomposition
In order to obtain decision components in the trajectory logs and selectively study them, we can decompose them through mathematical analysis tools. Fourier transform is a commonly used analysis method. It extracts the coefficient of the signal at different frequencies by convolution signal with sine and cosine functions as basis functions. However, since sine and cosine functions are infinite supporting, this decomposition will completely lose the time domain characteristics of the signal, resulting in our failing to select part of a certain time segment in each domain component. The wavelet transformation can help with this. Wavelets, also known as scale-varying basis functions, refer to a method of decomposing signals using compact supporting functions as basis functions. For example, the Haar wavelet, the simplest basis function of the wavelet transformation, enjoys widespread application in signal processing. The basis functions of Haar wavelet are
@@
where $\mathcal{T}[k]$ is the discrete trajectory profile, $\mathcal{T}_{At}$ is the low-frequency approximation coefficient at time index $t$, $\mathcal{T}_{Dt}$ is the high-frequency detail coefficient at time index $t$, $k$ can be considered as the convolution handler.
Figure 4: The iterative detail decoder’s architecture leverages reference lane encoding as the initial query for multi-layer attention mechanisms. Detail coefficients extracted during each iteration are integrated with the preceding query to formulate the subsequent query, thereby refining the extraction of further detailed information. The initial query serves as an approximation, with the first query and the ensuing generated details being processed by an MLP. These details are then mapped and compared against the detailed components at each level of the ground truth trajectory. The profile residual values are supervised by the profile details within the decision scope. Notably, the horizon masks can be configured individually across levels.
@@
we decompose trajectory profiles into $N$ levels. The approximation obtained in the last level is nominated as the preliminary decision. With extraction of a detail component in each level, the length of the data is reduced by half. An illustration of velocity decision scope diagram from log decomposition is provided in Fig. 3.
III-B Iterative Detail Decoder
Since the decomposed trajectory profile has been obtained, the question arises as to how it can be effectively utilized to guide the supervision of model training process. One potential method is to reconstruct the whole trajectory by truncated detail profiles. Nevertheless, this method requires assumption of data out of horizon to be zero, which introduces more deterministic prior in supervision signal. To facilitate the learning of detailed information, IDD is designed, which is capable of progressively generating details across multiple iterations. The architecture of the IDD is depicted in Fig. 4. In the decoding process, the encoding of reference lane serves as the query $Q_{0}$, and scenarios encoding $E_{sce}$ serves as key and value for a series of self-attention operations. The details extracted at each iteration are accumulated with the previous query, which is then utilized as the subsequent query. This iterative accumulation allows for the acquisition of increasingly refined trajectory profile. Mathematically, the process can be described as
$\tilde{Q}_{i-1}$ $=\text{SelfAttn}\left(Q_{i-1},\text{dim}=0\right),$
$\tilde{Q}_{i-1}$ $=\text{SelfAttn}\left(Q_{i-1}^{\prime},\text{dim}=1\right),$
$\Delta Q_{i}$ $=\text{CrossAttn}\left(\tilde{Q}_{i-1},E_{\text{sce}},E_{\text{sce}}\right),$
$Q_{i}$ $=\Delta Q_{i-1}+Q_{i-1}.$
In the decoding process, residual variables $\Delta Q_{i}$ generated during each iteration are recorded. Both $Q_{0}$ and logged $\Delta Q_{i}$ list are remapped by the same MLP decoder, i.e.,
$\tilde{\mathcal{T}}_{A}^{N}=\operatorname{MLP}\left(Q_{0}\right),\,\tilde{\mathcal{T}}_{D}^{l}=\operatorname{MLP}\left(\Delta Q_{N-l}\right).$

### v2

@@ Abstract
In the context of autonomous driving, learning-based methods have been promising for the development of planning modules. During the training process of planning modules, directly minimizing the discrepancy between expert-driving logs and planning output is widely deployed. In general, driving logs consist of suddenly appearing obstacles or swiftly changing traffic signals, which typically necessitate swift and nuanced adjustments in driving maneuvers. Concurrently, future trajectories of the vehicles exhibit their long-term decisions, such as adhering to a reference lane or circumventing stationary obstacles. Due to the unpredictable influence of future events in driving logs, reasoning bias could be naturally introduced to learning based planning modules, which leads to a possible degradation of driving performance. To address this issue, we identify the decisions and their corresponding time horizons, and characterize a so-called decision scope by retaining decisions within derivable horizons only, to mitigate the effect of irrational behaviors caused by unpredictable events. Several viable implementations have been proposed, among which batch normalization along the temporal dimension is particularly effective and achieves superior performance. It consistently outperforms baseline methods in terms of driving scores, as demonstrated through closed-loop evaluations on the nuPlan dataset. Essentially, this approach accommodates an appealing plug-and-play feature to enhance the closed-loop performance of other learning-based planning models.
The source code and associated videos are available at https://github.com/Rex-sys-hk/PlanScope.
@@ I Introduction
Learning-based autonomous driving systems have been a promising approach for the popularity of unmanned vehicles. In particular, learning from driving logs of experts is broadly adopted to supervise neural networks for the planning task. During the training process, distance error is commonly employed, which is to measure the discrepancy between the planned state sequence and the expert demonstrations, thereby optimizing parameters of the neural network [1]. Yet, when training end-to-end autonomous driving systems, it is not entirely reasonable to set the logged expert states as target sequence, because future events are inherently uncertain, and potential disruptions such as unexpected obstacles or fluctuating traffic signals can significantly alter the trajectory planning. These unpredictable elements necessitate agile and nuanced adjustments to driving behavior. Some seemingly predictable decisions, such as behaviors in Game models with multiple Nash Equilibrium and switching traffic light, are difficult to accurately predict. Meanwhile, a short-sighted planner is also not desirable, as the planner also needs to keep its long-term decision in consideration when choosing sudden maneuvers. Essentially, the state sequence of experts recorded in the log is a combination of short-term and long-term decisions under different scopes. A diagram of this concept is shown in Fig. 1.
@@
Due to the lack of observable information, long-term decisions are more susceptible to unpredictable events compared to short-term decisions. In learning targets like Average Distance Error (ADE), waypoints near the given current state only account for a small proportion of error compared to remote ones. This will result in a situation where unpredictable events occurring at the distant future affect current short-term decisions unexpectedly. This issue is evident during the training phase of the planning model, where it can lead to uneven convergence priorities or errant convergence over time. Therefore, considering the decision scope (DS) in the model training phase is crucial for making more intelligent planning decisions.
Addressing the decoupling issue of information from various decision scopes within the learning-from-demonstration framework presents significant challenges. Traditional planning methods divide the problem into distinct layers: path planning, behavior planning, and motion planning. Each of these layers is handled by specialized modules. In contrast, learning-from-demonstration approaches often focus on learning future trajectories of a fixed length. This focus arises because nuanced adjustment components and long-term decision components are difficult to separate from log replay. One potential solution is to employ manual annotation for driving behavior reasoning. This method can help analyze both long-term and short-term decisions, allowing for the identification of behavior duration and starting times. However, this approach entails a substantial workload and poses scalability challenges. Consequently, finding a cost-effective way to decouple decision scopes in the age of learning-based planning remains a complex issue that requires further exploration.
Three viable strategies are investigated in this work: 1) Assigning weight to the time dimension of loss; 2) Monitoring tracks through decisions of varying scales; and 3) Enabling the model to refine trajectories from coarse to fine. In our examination of time-dependent weighting methods, three approaches are proposed: time truncation, uncertainty-dependent weighting, and time-dependent normalization. These weights are incorporated into the computation of planning regression loss. For monitoring decision scales, we suggest that the model can learn to make decisions across different scopes by employing downsampling and truncation. Alternatively, the model can extract components of varying frequencies from expert trajectories using wavelet decomposition and adjust them according to planning requirements. Regarding the model structure for generating decisions at different scopes, both one-time and iterative generation methods are explored, which enable the model to refine trajectories from coarse to fine. To ensure a comprehensive and effective exploration, experimental validations of these strategies are conducted. Through our experiments, the most effective method is identified for enhancing model performance in closed-loop evaluations.
In summary, our contributions are threefold:
  • We introduce PlanScope, a framework for formulating problems on training the sequential decision-making model, and demonstrate its impact by closed-loop evaluations.
  • We explore a variety of methods under PlanScope framework in the aspects of learning weight calculation, trajectory decomposition, and the modeling structure of detail generation. Through comparative experiments, the most effective strategies are identified for enhancing trajectory execution performance.
@@
  • We extensively evaluate our method on the nuPlan [2] dataset, and the results demonstrate that it achieves superior driving scores compared to baseline methods in closed-loop none-reactive simulations (CLS-NR).
@@ II-A Horizon Related Planner Design
Sequential decision-making can be encapsulated with the framework of Partially Observable Markov Decision Processes (POMDPs) [3]. For tasks with fixed decision horizons, this framework requires exponentially increasing search volume to find a feasible solution. However, in autonomous driving scenarios, in-time reaction and predictive decision are both required, making it challenging to determine the optimal policy sequence in real-time [4]. Hence, under traditional planning frameworks, hierarchical structure is typically employed to streamline computations [5, 6]. In the hierarchical structure, the planning problem is decomposed across various frequency bands, each with distinct hierarchical tendencies. For large-scale exploration, one might opt for a lower resolution and focus on non-real-time targets. Conversely, within the smaller scale areas identified by the large-scale search, the planner can leverage higher resolution and account for dynamic objects that are impractical to predict accurately. In [7], the authors adapts this concept to the prediction phase, and segments the forecast for the subsequent four seconds into two stages. Initially, it meticulously predicts the agent’s reachable area at 0.1-second intervals for the first 0.5 seconds. Subsequently, for the remaining 3.5 seconds, it utilizes an adaptive filter to depict potential vehicle positions. This approach allows the agent to maintain safety at a fine scale while preserving adaptability at a broader scale. In terms of learning-based planning, analogous strategies can be harnessed to simplify calculations and enhance performance. For instance, the Negative Log Likelihood (NLL) Loss, which is frequently utilized in predictive tasks [8], assigns greater uncertainty to long-term trajectory points, such that the scarcity of future information can be reflected. This phenomenon of probabilistic regression result helps to posit that historical trajectories provide limited insight, sufficient only for fine-tuned operations in the short term and for general directional guidance in the long term. However, NLL loss or probabilistic predictions are incompatible with the planning module, which necessitates deterministic outputs for trajectory followers. It leaves an open and interesting problem to develop an approach to address this challenge in trajectory planning. Essentially, there is a critical need for a method that accurately reflects the significance of decision components in trajectories or can extract them effectively.
@@ III Methodology
Figure 2: The overall model framework initiates with the individual embedding of driving context elements, including the Ego Vehicle, Dynamic Agents, Static Objects, High-Definition Map, Traffic Signs, and Signals, followed by their concatenation with positional embedding (PE) [15]. This concatenated context is subsequently processed through a Multi-layer Transformer encoder, iterated $L_{enc}$ times. The scenario embedding is then decoded to formulate agent predictions, which serve to comprehensively supervise the model’s understanding of the driving context. The reference lane embedding and anchor-free variables are combined to form a query for the scenario embedding.
III-A Problem Formulation & System Overview
@@
where $f$ denotes the Neural Network (NN) model, $\theta$ is the model parameter, $\mathcal{L}$ is the loss function, and $\hat{G}_{t}$ is the ground truth of $G_{t}$ from logged context.
@@
where $\mathcal{L}_{\text{pre}}$ calculates the future trajectory discrepancy between generated and that from dataset, $\mathcal{L}_{\text{col}}$ computes the invasion distance using $N_{c}$ circle represented vehicle outline in future $T_{f}$ steps as described in [19], $R_{c}$ is the summed radius of ego and agents outline circle, $\epsilon$ is the tolerance, and $d_{i}^{t}$ represents the center distance between outline circles at time step $t$. Drawing inspiration from [20, 21, 22], the decoder’s query is formulated as a concatenation of the reference lane encoding and anchor-free variables. The reference lane encoded features are concatenated with anchor-free learnable queries and serve as the initial query $Q_{0}$ to the scenario encoding within multi-layer tranformer decoding operations, facilitating the generation of residual quantities of $Q$ for subsequent layers. In training process, the final query $Q_{N}$ is then forwarded to the trajectory MLP decoder and scoring MLP decoder, yielding the multi-modal trajectory and the score by
$\mathcal{T}$ $=\operatorname{MLP}(Q_{N}),\,\pi=\operatorname{MLP}(Q_{N}),$
$\mathcal{L}_{\text{cls}}$ $=\operatorname{CrossEntropy}\left(\pi,\hat{\pi}\right),$
$\mathcal{L}=\mathcal{L}_{\text{reg}}+\mathcal{L}_{\text{cls}}+\mathcal{L}_{\text{pre}}+\mathcal{L}_{\text{col}}+\mathcal{L}_{\text{ds}},$
where $\mathcal{L}_{\text{reg}}$ is the weighted regression loss and $\mathcal{L}_{\text{ds}}$ is the loss to supervise learning within decision scope. More details on weights design, acquisition of supervision signals and the decoding of detailed information are elaborated in Section III-B, Section III-C and Section III-D2, respectively.
III-B Weighted Supervision
Figure 3: Frameworks for achieving multi-scope trajectory supervision are illustrated in (a), (b), and (c). In (a), the trajectory is generated using the baseline framework but is supervised by a loss function with time-dependent weights. In (b), the last trajectory embedding is used not only to generate the full trajectory but also to map to the decomposed trajectory. In (c), multi-scope detail embeddings are iteratively generated during the decoding process of the Transformer Decoder.
Using time-dependent weight along decision sequence is a promising way to express the importance of waypoint accuracy in the learning process as shown in Fig. 3 (a). The weighted regression loss can be calculated by
$\begin{array}[]{c}\mathcal{L}_{\text{reg}}{}_{t}=\operatorname{L1_{smooth}}\left(\mathcal{T}_{t},\hat{\mathcal{T}}_{t}\right),\\ \mathcal{L}_{\text{reg}}=\frac{1}{T}\sum^{T}_{t=0}\mathcal{L}_{\text{reg}}{}_{t}\times w_{t},\end{array}$
where $w_{t}$ can be designated by the following methods:
III-B1 Time Truncation
Only consider regression loss of first T time steps. In implementation, we also shorten the executed trajectory to eliminate the effect of uncontrolled waypoints. Mathematically,
$w_{t}=\left\{\begin{array}[]{ll}1,&t<T,\\ 0,&t\geq T.\end{array}\right.$
III-B2 Decreasing Certainty
Those unpredictable disturbances might affect convergence capability on more important features and information for immediate decision. To describe this problem, we assume the sequential decision process conforms to the Gaussian Process (GP). GP is a non-parametric Bayesian method used to model functions. It assumes that any finite collection of function values follows a joint Gaussian distribution. For trajectory planning of an autonomous vehicle, the trajectory points $\mathbf{y}(t)$ can be modeled as outputs of a GP, with the mean and covariance determined by a kernel function $k(t_{i},t_{j})$, which can be a Radial Basis Function (RBF). Assuming the current state is an observable point, the variance of the predicted point at $t$ can be easily derived by
$\sigma_{t}^{2}=\sigma_{f}^{2}\left[1-\exp\left(-\frac{(t-t_{0})^{2}}{l^{2}}\right)\right],$
where $\sigma_{f}^{2}$ is the function uncertainty, which can also be considered as the maximum uncertainty, $l$ is the time length scaler, $t_{0}$ is the observable current time, which is set to 00 in the calculation. We define the uncertainty compensation as the deviation of maximum uncertainty. Mathematically,
$\text{Compensation}(t)=\sigma_{f}^{2}-\sigma_{t}^{2}.$
Then, the calculation of weights to regularize uncertainty at each time point and can be simplified as
$w_{t}=\frac{1}{Z}\operatorname{exp}(-(\frac{t-t_{0}}{l})^{p}),$
where $Z$ is the average weight. $p$ is the order of the time difference measurement, which is $2$ for RBF.
III-B3 Time Dependent Normalization
Normalization is a straight forward method to convert value of data to the same order of magnitude or limits it to a certain range. Borrowing the idea of Batch Normalization (BN), the average error $e_{t}$ of each time step across the mini-batch can be dynamically calculated by
$BN(e_{t})=\frac{e_{t}-\mu_{B;t}^{e_{t}}}{\sigma_{B;t}}=\frac{e_{t}}{\mu_{B;t}^{|e_{t}|}},\text{ when }\mu_{B;t}^{e_{t}}=0.$
So distance error at each timestep $e_{t}$ can be normalized by multiplying with weight
$\begin{array}[]{c}w_{t}=\left(\frac{1}{B}\sum^{B}_{b=1}\mathcal{L}_{reg}{}_{b;t}\right)^{-1},\end{array}$
where $b$ is the index of data item in mini-batch and $B$ is the batch size.
III-C Expert Log Decomposition
@@
In order to obtain decision components in the trajectory logs and selectively study them, we can decompose them through mathematical analysis tools. Fourier transform is a commonly used analysis method. It extracts the coefficient of the signal at different frequencies by convolution signal with sine and cosine functions as basis functions. However, since sine and cosine functions are infinitely supportive, this decomposition will completely lose the time domain characteristics of the signal, resulting in our failure to select part of a certain time segment in each domain component. To achieve time dependent decomposition of trajectories, we propose the following two approaches.
III-C1 Downsampling within Horizon (DWH)
In this method, we do not utilize any signal feature extracting operation. But only downsample and truncate the limited horizon of future time steps of expert trajectory iteratively. The downsampling and truncation operations are adopted N times to be aligned with details generated by the NN. Mathematically,
$\begin{array}[]{c}\mathcal{T}_{l*}=\mathcal{T}\downarrow 2^{l-1},\\ \mathcal{T}_{l}=\mathcal{T}_{l*}{}_{1:H},\end{array}$
where $H$ is the Horizon of decision in each detailing level $l$.
III-C2 Wavelet Decomposition
Wavelets, also known as scale-varying basis functions, refer to a method of decomposing signals using compact supporting functions as basis functions. For example, the Haar wavelet, the simplest basis function of the wavelet transformation, enjoys widespread application in signal processing. The basic functions of the Haar wavelet are
@@ III-C2 Wavelet Decomposition
where $\mathcal{T}[k]$ is the discrete trajectory profile, $\mathcal{T}_{At}$ is the low-frequency approximation coefficient at time index $t$, $\mathcal{T}_{Dt}$ is the high-frequency detail coefficient at time index $t$, $k$ can be considered as the convolution handler. For Haar wavelet decomposition, $\mathcal{T}_{At}$ represents the sum of $\mathcal{T}[k]$ and $\mathcal{T}[k+1]$ while $\mathcal{T}_{Dt}$ is the product similarities bwtween $[\mathcal{T}[k],\mathcal{T}[k+1]]$ and $[1,-1]$.
Figure 4: The iterative detail decoder’s architecture leverages reference lane encoding as the initial query for multi-layer attention mechanisms. Detail coefficients extracted during each iteration are integrated with the preceding query to formulate the subsequent query, thereby refining the extraction of further detailed information. The initial query serves as an approximation, with the first query and the ensuing generated details being processed by an MLP. These detail embeddings are then mapped and compared against the detailed components at each level of details. Notably, the horizon masks can be configured individually across levels.
@@
we decompose trajectory profiles into $N$ levels. The profile in our experiments is selected to be coordinate values of waypoints. The approximation obtained in the last level is nominated as the preliminary decision. With the extraction of a detail component in each level, the length of the data is reduced by half.
III-D Detail Generation
Since the decomposed trajectory profile has been obtained, the question arises as to how it can be effectively utilized to guide the supervision of the model training process.
III-D1 Multi-head Detail Decoder (MDD)
This method leverages the function-fitting ability of MLP to keep gradients. As shown in Fig. 3 (b), the transformer decoder layer generates trajectory embedding. MLPs are utilized to remap trajectory embedding to details of designated levels and the overall approximation. Mathematically,
$\tilde{\mathcal{T}}_{A}^{N}=\operatorname{MLP}\left(Q_{N}\right),\,\tilde{\mathcal{T}}_{D}^{l}=\operatorname{MLP}_{l}\left(Q_{N}\right).$
III-D2 Iterative Detail Decoder (IDD)
To facilitate the learning of detailed information, IDD is designed, which is capable of progressively generating details across multiple iterations as shown in Fig. 3 (c). The architecture of the IDD is depicted in Fig. 4. In the decoding process, the embedding of the reference lane serves as the query $Q_{0}$, and scenarios encoding $E_{sce}$ serves as the key and value for a series of self-attention operations. We hypothesize that the multi-layer transformer decoder as a trajectory refining process. This hypothesis is based on the observation that the layer normalization operation within each decoding layer reduces information capacity while enhancing accuracy. The embedding extracted at each iteration are accumulated with the previous query, which is then utilized as the subsequent query. Mathematically, the process can be described as
[…]
In the decoding process, the transformed feature $Q_{i}$ is generated during each iteration and recorded. Logged $\{Q_{i}\}$ is adapted by level-wise MLP decoders, i.e.,
$\tilde{\mathcal{T}}_{A}^{N}=\operatorname{MLP}_{A}\left(Q_{0}\right),\,\tilde{\mathcal{T}}_{D}^{l}=\operatorname{MLP}_{l}\left(Q_{l}\right).$